# Gradient Descent with Large Step Sizes: Chaos and Fractal Convergence Region

### v2

@@ Abstract
We examine gradient descent in matrix factorization and show that under large step sizes the parameter space develops a fractal structure. We derive the exact critical step size for convergence in scalar-vector factorization and show that near criticality the selected minimizer depends sensitively on the initialization. Moreover, we show that adding regularization amplifies this sensitivity, generating a fractal boundary between initializations that converge and those that diverge. The analysis extends to general matrix factorization with orthogonal initialization. Our findings reveal that near-critical step sizes induce a chaotic regime of gradient descent where the long-term dynamics are unpredictable and there are no simple implicit biases, such as towards balancedness, minimum norm, or flatness.
@@ 1 Introduction
Understanding the properties of gradient descent in non-convex overparametrized optimization has been a central pursuit in modern machine learning. The step size, or learning rate, is a critical factor determining the dynamics and convergence of gradient descent optimization. In particular, it has a major influence on the returned solution and its generalization performance (Nar and Sastry,, 2018; Jastrzebski et al.,, 2020; Lewkowycz et al.,, 2020; Cohen et al.,, 2021). Large step sizes have been associated with flat and balanced minimizers of the training objective (Wu et al.,, 2018; Wang et al.,, 2022; Menon,, 2024), sparse feature representations (Nacson et al.,, 2022; Andriushchenko et al.,, 2023), smooth solution functions (Mulayoff and Michaeli,, 2020; Nacson et al.,, 2023), and improved generalization (Ba et al.,, 2022; Qiao et al.,, 2024; Sadrtdinov et al.,, 2024). Yet, the theoretical understanding of large step sizes remains limited, even in simple convex settings. Our investigation is motivated by two fundamental questions:
@@
Addressing these questions is challenging, since large step sizes can produce highly complex, non-monotonic, and even chaotic trajectories. In particular, trajectories may not converge to stationary points but instead enter periodic or chaotic oscillations (Chen and Bruna,, 2023; Chen et al., 2024b, ; Ghosh et al.,, 2025), or converge to a statistical distribution (Kong and Tao,, 2020); trajectories that eventually converge to a minimizer may still undergo chaotic oscillations during early training (Zhu et al.,, 2023; Kreisler et al.,, 2023; Song and Yun,, 2023); and trajectories with nearby initializations can diverge exponentially from one another (Herrmann et al.,, 2022; Jiménez-González et al.,, 2025). Moreover, empirically, the set of step sizes and the set of initializations leading to convergence can form fractal structures (see, respectively, Sohl-Dickstein,, 2024; Zhu et al.,, 2023). In this work, we provide precise answers to the above questions in the context of matrix factorization problems with rigorous theoretical characterizations.
@@
We begin by examining gradient descent in a simplified problem to factor a scalar target as the inner product of two vectors. We show that two striking phenomena emerge at large step sizes: (i) trajectories originating from arbitrarily small sets of initializations can converge to global minimizers of the training loss with arbitrarily large norm, sharpness or imbalance, or to a saddle point; and (ii) the set of initializations that converge, that is, the convergence region, has a fractal structure (see the left panel in Figure 1). Thus gradient descent exhibits sensitivity to initialization and its long-term behavior is unpredictable. Interestingly, while the convergence region is regular (in the almost everywhere sense) in the unregularized setting, it becomes fractal once regularization is introduced. Further quantifying the unpredictability of gradient descent, we show that the topological entropy of the gradient descent system is at least $\log 3$. Also, we show that the fractal nature of the boundary of the convergence region is captured by a self-similar curve, whose fractal dimension is estimated as $1.249$. To our knowledge, beyond the univariate training loss setting studied by Kong and Tao, (2020); Chen et al., 2024b , this is the first rigorous characterization of chaos in gradient descent optimization.
@@
Figure 1: Left: Gradient descent applied to $L(u,v)=(u^{\top}v-1)^{2}+0.3(\|u\|_{2}^{2}+\|v\|_{2}^{2})$ with $(u,v)\in\mathbb{R}^{10}$. Shown is a random two-dimensional slice of $\mathbb{R}^{10}$. Gray points are initializations from which the algorithm does not converge; other points are colored by the value of one of the coordinates of the converged minimizer. As we see, the convergence boundary is fractal, and the converged solution depends sensitively on the initialization when this is near the boundary. Right: Gradient descent applied to $L(x,y)=(xy-1)^{2}$ with $(x,y)\in\mathbb{R}^{2}$. The green star marks a balanced minimizer $p$, with its neighborhood $O_{p}$ depicted as a blue disk with black boundary. The blue region with dashed boundaries shows the preimage of $O_{p}$ under $\mathrm{GD}^{6}$. The green diamond marks an imbalanced minimizer, with its neighborhood and preimages shown in orange. For this problem the convergence boundary is smooth but the convergence point for initializations near the boundary is chaotic.
@@
We then extend our analysis to general matrix factorization by showing that, when the initialization lies in a subspace defined by a set of orthogonal conditions, the gradient descent dynamics decouples into several independent scalar factorization dynamics. Hence, all the results established for scalar factorization remain valid and can be applied on this subspace. This includes, in particular, the commonly used identity initialization. Our results also extend to shallow linear residual networks (Hardt and Ma,, 2017; Bartlett et al.,, 2018). We experimentally show that sensitivity to initialization persist over more general settings, such as generic initialization and deep matrix factorization.
We further analyze the mechanisms underlying these phenomena and trace them to a folding behavior of the update map $\mathrm{GD}(\theta)=\theta-\eta\nabla L(\theta)$: the map $\mathrm{GD}$ sends a region $\mathcal{C}$ containing multiple minimizers onto a larger region that contains $\mathcal{C}$, in a multi-fold covering manner. As a consequence, for any neighborhood $O_{p}$ of a minimizer $p\in\mathcal{C}$, the number of connected components in the preimage $\mathrm{GD}^{-N}(O_{p})$} grows exponentially with the number of iterations $N$, and these components accumulate near the boundary of the convergence region. The preimages associated with different minimizers then become intricately interwoven near the boundary, giving rise to sensitive dependence on the initialization and a self-similar fractal structure (see the right panel in Figure 1). Overall, our results show that near-critical step sizes place gradient descent in a chaotic regime, where infinitesimal perturbations on the initial conditions can lead to substantially different training outcomes. This stands in sharp contrast to the stable dynamics observed at smaller step sizes.
1.1 Main Contribution
@@
  • We derive the exact critical step size for convergence in scalar factorization and show that, at critical step size, infinitesimal perturbations of the initialization can redirect gradient descent to global minimizers with arbitrarily large norm, imbalance, or sharpness, or to a saddle. We show that the topological entropy of gradient descent system is at least $\log 3$. We also show that the convergence region in the parameter space is equal almost everywhere to a bounded, smooth domain.
  • We show that, with $\ell_{2}$ regularization, gradient descent selects either the minimal distance solution or the maximal distance solution among all global minimizers. Near criticality, infinitesimal perturbations of the initialization can switch this selection from one to the other. We show that if the step size is below an explicit bound, the algorithm always selects the minimal distance solution.
  • We show that adding $\ell_{2}$ regularization yields a fractal convergence boundary, whose geometry is captured by a self-similar shape in $\mathbb{R}^{2}$ after symmetry reduction. We numerically estimate the fractal dimension of this shape. We further show that, up to a measure-zero set, the convergence region has an unbounded interior.
  • We extend these results to general matrix factorization under a range of initializations including the identity initialization. We show that gradient descent partitions the parameter space into components where it acts as a covering map, and reason that chaos arises when it is an expansion on a component that contains multiple minimizers.
@@ 1.2 Related Work
A main line of research on large-step-size gradient descent focuses on the non-monotonic convergence of the loss and its impact on the final model. Key perspectives include the Edge of Stability (Cohen et al.,, 2021; Ma et al.,, 2022; Agarwala et al.,, 2023; Damian et al.,, 2023; Ahn et al.,, 2022, 2023; Zhu et al.,, 2023; Wang et al.,, 2023) and the catapult phenomenon (Lewkowycz et al.,, 2020; Kalra and Barkeshli,, 2023; Meltzer and Liu,, 2023; Zhu et al., 2024a, ; Zhu et al., 2024b, ). Compared to these works, our analysis extend to even larger (near-critical) step sizes. Another line of work shows how a large step size can enhance feature learning in one step of gradient descent (Ba et al.,, 2022; Dandi et al.,, 2024; Moniri et al.,, 2025), also comparing different parametrizations (Sonthalia et al.,, 2025). Similar observations about the role of the step size in feature learning have also been made in SGD (Andriushchenko et al.,, 2023; Lu et al.,, 2024) and pre-training (Sadrtdinov et al.,, 2024). Ziyin et al., (2022) observed that for a certain range of step sizes, SGD can have undesirable behavior, such as convergence to local maxima. For linear networks, Kreisler et al., (2023) identified a monotonically decreasing quantity (sharpness) along the gradient descent trajectories. Wang et al., (2022) showed large step size induces an implicit bias towards balanced minimizers in matrix factorization. Crăciun and Ghoshdastidar, (2024) proved the existence of a step size threshold above which the algorithm diverges. Large-step-size gradient descent has also been investigated in logistic regression (Wu and Su,, 2023; Wu et al.,, 2024; Meng et al.,, 2024), and some of the analysis has been further extended to shallow networks (Cai et al.,, 2024).
@@
Van Den Doel and Ascher, (2012) empirically observed chaos, specifically positive finite-time Lyapunov exponents, for several variants of steepest descent methods. The phenomenon named period-doubling bifurcation route to chaos has been widely observed in recent literature (Kong and Tao,, 2020; Chen and Bruna,, 2023; Chen et al., 2024b, ; Meng et al.,, 2024; Danovski et al.,, 2024; Ghosh et al.,, 2025). Among them, only Kong and Tao, (2020); Chen et al., 2024b provided rigorous analyses for the chaotic dynamics. They showed the emergence of Li-Yorke chaos, i.e., the existence of periodic orbits of arbitrary periods, for univariate training losses. In comparison, our setting is high-dimensional. Additionally, we established not only the existence of all periodic orbits, but also the sensitivity of the limiting point to initialization, which is more relevant to practical optimization, particularly the implicit bias of the optimization algorithm.
@@ 2 Preliminaries
Gradient optimization in problem (1) has been extensively studied, especially in small-step-size regimes (see, e.g., Saxe et al.,, 2013; Arora et al.,, 2019; Yun et al.,, 2021; Du et al.,, 2018; Li et al.,, 2021; Min et al.,, 2023; Chen et al., 2024a, ). Its landscape enjoys a favorable structure: the global minimum is attained and every stationary point is either a global minimizer or a strict saddle (Li et al., 2019b, ; Valavi et al.,, 2020; Zhou et al.,, 2022). Nevertheless, this problem retains a key complexity typical of neural network optimization: global minimizers can differ substantially (although they yield the same end-to-end matrix). In particular, in the unregularized case, both the parameter norm $\|U\|_{F}^{2}+\|V\|_{F}^{2}$ and the imbalance $\|UU^{\top}-VV^{\top}\|_{F}^{2}$ can be arbitrarily large on the set of minimizers. This makes the problem a natural testbed for studying how hyperparameter choices affect the implicit biases of parameter optimization algorithms. We note that other forms of regularization have also been studied in the literature, such as $\|UU^{\top}-VV^{\top}\|_{F}^{2}$ (Tu et al.,, 2016; Ge et al.,, 2017).
@@
We define the gradient descent update map $\mathrm{GD}_{\eta}(\theta)=\theta-\eta\nabla L(\theta)$ so that $(U_{t+1},V_{t+1})=\mathrm{GD}_{\eta}(U_{t},V_{t})$. The basin of attraction of a stationary point $\theta^{*}$ of $L$ is the set of all initializations that converge to $\theta^{*}$, $\{\theta\colon\lim_{N\to\infty}\mathrm{GD}^{N}_{\eta}(\theta)=\theta^{*}\}$  [^1] In dynamical systems the basin of attraction is often defined for attractors. Here we extend the terminology to include all stationary points, such as saddles, for simplicity of presentation.. The convergence region for step size $\eta$, denoted by $\mathcal{D}_{\eta}$, is the union of the basins of attraction of all global minimizers. The critical step size $\eta(\bar{U},\bar{V})$ for an initialization $(\bar{U},\bar{V})$ is defined as $\eta(\bar{U},\bar{V})=\sup\left\{\eta\colon\lim_{N\to\infty}\mathrm{GD}^{N}_{\eta}(\theta)\in\mathcal{M}\right\}$, i.e., the supremum of the step sizes that allow convergence, where $\mathcal{M}$ denotes the set of all global minimizers.
@@
We introduce notions for describing fractal geometry. A fractal is typically defined as a shape that exhibits self-similarity and fine structure at arbitrarily small scales. Formally, we say a set $S\subset\mathbb{R}^{n}$ is self-similar with degree k if there exist $k$ homeomorphisms, $\phi_{i}\colon S\to S,i=1,\cdots,k$, that satisfy (i) $S=\cup_{i=1}^{k}\phi_{i}(S)$ and (ii) there exists an open set $O\subset S$ such that $\cup_{i=1}^{k}\phi_{i}(O)\subset O$ and $(\phi_{i}(O))_{i=1}^{k}$ are pairwise disjoint. Condition (i) states that $S$ can be covered by $k$ smaller copies of itself. Condition (ii), which is known as the open set condition, ensures that those copies do not overlap much. This definition is closely related to an Iterated Function System (IFS), a standard tool for analyzing fractals (see, e.g., Hutchinson,, 1981; Falconer,, 2013). However, unlike IFS where the maps are required to be contractive and the set $S$ to be compact, the shapes considered in our study may be unbounded.
@@
Finally, we introduce notions related to chaos. Although there is no universal definition of chaos, one common characterization of chaos is the sensitivity to initialization, which is often known as the butterfly effect. In the context of optimization, this manifests as the phenomenon where infinitesimal perturbations of the parameter initialization or step size can lead to substantially different training outcomes (e.g., turning convergence into divergence, or shifting convergence from one minimizer to another qualitatively different minimizer). In dynamical systems, one of the most important measures of chaos is the topological entropy. Informally, the topological entropy $h(F)$ of a dynamical system $F$ measures the exponential growth rate of the number of distinct trajectories of $F$ as a function of the trajectory length. We defer the formal definition to Appendix B.1. A positive topological entropy is widely regarded as a hallmark of chaos (see, e.g., Katok et al.,, 1995; Robinson,, 1998; Vries,, 2014). In this paper, we adopt the above notions for fractals and chaos. We note that different definitions and settings exist, and discuss their relation to our study in Appendix A.

### v3

@@ Abstract
We examine gradient descent in matrix factorization and show that under large step sizes the parameter space develops a fractal structure. We derive the exact critical step size for convergence in scalar-vector factorization and show that near criticality the selected minimizer depends sensitively on the initialization. Moreover, we show that adding regularization amplifies this sensitivity, generating a fractal boundary between initializations that converge and those that diverge. The analysis extends to general matrix factorization with orthogonal initialization. Our findings reveal that near-critical step sizes induce a chaotic regime of gradient descent where the training outcome is unpredictable and there are no simple implicit biases, such as towards balancedness, minimum norm, or flatness.
@@ 1 Introduction
Understanding the properties of gradient descent in non-convex overparametrized optimization has been a central pursuit in modern machine learning. The step size, or learning rate, is a critical factor determining the dynamics and convergence of gradient descent optimization. In particular, it has a major influence on the returned solution and its generalization performance (Nar and Sastry, 2018; Jastrzebski et al., 2020; Lewkowycz et al., 2020; Cohen et al., 2021). Large step sizes have been associated with flat and balanced minimizers of the training objective (Wu et al., 2018; Wang et al., 2022; Menon, 2025), sparse feature representations (Nacson et al., 2022; Andriushchenko et al., 2023), smooth solution functions (Mulayoff and Michaeli, 2020; Nacson et al., 2023), and improved generalization (Ba et al., 2022; Qiao et al., 2024; Sadrtdinov et al., 2024). Yet, the theoretical understanding of large step sizes remains limited, even in simple convex settings. Our investigation is motivated by two fundamental questions:
@@
Addressing these questions is challenging, since large step sizes can produce highly complex, non-monotonic, and even chaotic trajectories. In particular, trajectories may not converge to stationary points but instead enter periodic or chaotic oscillations (Chen and Bruna, 2023; Chen et al., 2024b; Ghosh et al., 2025), or converge to a statistical distribution (Kong and Tao, 2020); trajectories that eventually converge to a minimizer may still undergo chaotic oscillations during early training (Zhu et al., 2023; Kreisler et al., 2023; Song and Yun, 2023); and trajectories with nearby initializations can diverge exponentially from one another (Herrmann et al., 2022; Jiménez-González et al., 2025). Moreover, empirically, the set of step sizes and the set of initializations leading to convergence can form fractal structures (see, respectively, Sohl-Dickstein, 2024; Zhu et al., 2023). In this work, we provide precise answers to the above questions in the context of matrix factorization problems with rigorous theoretical characterizations.
@@
We begin by examining gradient descent in a simplified problem to factor a scalar target as the inner product of two vectors. We show that two striking phenomena emerge at near-critical step sizes: (i) initializations in arbitrarily small sets can converge to global minimizers with arbitrarily large norm, sharpness or imbalance, or to a saddle, indicating a lack of simple implicit biases; and (ii) the set of initializations that converge, that is, the convergence region, has a fractal structure, indicating that the convergence is unpredictable near the boundary (see Figure 1). Interestingly, while in the unregularized setting the convergence region is regular (in the almost everywhere sense), it becomes fractal once regularization is introduced. Further quantifying the unpredictability of the training outcome, we show that the topological entropy of the gradient descent system is at least $\log 3$. Also, we show that the fractal nature of the boundary of the convergence region is captured by a self-similar curve whose fractal dimension is estimated as $1.249$. To our knowledge, beyond the univariate training loss setting studied by Kong and Tao (2020); Chen et al. (2024b), this is the first rigorous characterization of chaos in gradient descent optimization.
@@
Figure 1: The training outcome of gradient descent depends sensitively on initializations near the convergence boundary. Left: Gradient descent applied to $L(u,v)=(u^{\top}v-1)^{2}+0.3(\|u\|_{2}^{2}+\|v\|_{2}^{2})$ with $(u,v)\in\mathbb{R}^{10}$. Shown is a random two-dimensional slice of $\mathbb{R}^{10}$. Gray points are initializations leading to divergence; other points are colored by a coordinate value of the converged minimizer. The convergence boundary has a fractal structure. Right: Gradient descent applied to $L(u,v)=(uv-1)^{2}$ with $(u,v)\in\mathbb{R}^{2}$. Shown are a balanced minimizer $p$ (green star), its neighborhood $O_{p}$ (blue disk), and the preimage of $O_{p}$ under $\mathrm{GD}^{6}$ (blue region with dashed boundary). An imbalanced minimizer (green diamond) is shown similarly, with its neighborhood and preimages depicted in orange. The convergence boundary is smooth but the converged point is sensitive to initialization.
@@
We then extend our analysis to general matrix factorization by showing that, when the initialization lies in a subspace defined by a set of orthogonal conditions, the gradient descent dynamics decouples into several independent scalar factorization dynamics. Hence, all results established for scalar factorization remain valid on this subspace. This includes the commonly used identity initialization, as well as the training of linear residual networks (Hardt and Ma, 2017; Bartlett et al., 2018).
We further analyze the mechanisms underlying these phenomena and trace them to a folding behavior of the update map $\mathrm{GD}(\theta)=\theta-\eta\nabla L(\theta)$: the map $\mathrm{GD}$ sends a region $\mathcal{C}$ onto a superset of $\mathcal{C}$ in a multi-fold covering manner. Consequently, if $\mathcal{C}$ contains a convergence boundary that is invariant under $\mathrm{GD}$, then the boundary exhibits self-similarity; moreover, $\mathrm{GD}$ is mixing on the boundary, giving rise to the sensitivity to initialization. We show that for general neural networks with polynomial activations $\mathrm{GD}$ indeed acts as a covering map on the parameter space outside a measure-zero set.
Overall, our results imply that near-critical step sizes place gradient descent in a chaotic regime where the training outcome is unpredictable: infinitesimal perturbations on the initial conditions can lead to substantially different outcomes. This contrasts sharply with the stable dynamics observed at smaller step sizes. We empirically validate the presence of chaos in matrix factorization with general initializations, deep matrix factorization, and deep ReLU networks trained on real-world datasets.
1.1 Main Contributions
@@
  • We derive the exact critical step size for convergence in scalar factorization and show that the convergence region is equal almost everywhere to a bounded and smooth domain. At critical step sizes, we show that infinitesimal perturbations to the initialization can lead to global minimizers with arbitrarily large norm, or to a saddle point. Moreover, for initializations randomly sampled from arbitrarily small sets, the distribution of the sharpness at the converged minimizers has a support containing all possible values below $2/\eta$, which evidences chaos at a distributional level. Also, we show that the topological entropy of gradient descent system is at least $\log 3$.
  • We show that when using $\ell_{2}$ regularization, gradient descent selects either the minimal distance solution or the maximal distance solution among all global minimizers. At critical step size, infinitesimal perturbations of the initialization can switch this selection from one to the other. Furthermore, adding $\ell_{2}$ regularization yields a fractal convergence boundary whose geometry is captured by a self-similar shape in $\mathbb{R}^{2}$ after symmetry reduction.
  • We extend these results to general matrix factorization with initializations in a subspace that includes the identity initialization. In particular, gradient descent exhibits chaos and fractal convergence boundary on this subspace.
  • We reason that chaos arises from a folding behavior of the gradient descent update map. For general neural networks with polynomial activations, we show that gradient descent is a covering map on each connected component of the parameter space after removing a measure-zero set. We empirically validate these chaotic phenomena in deep ReLU networks trained on real-world data.
@@ 1.2 Related Work
A main line of research on large-step-size gradient descent focuses on the non-monotonic convergence of the loss and its impact on the final model. Key perspectives include the Edge of Stability (Cohen et al., 2021; Ma et al., 2022; Agarwala et al., 2023; Damian et al., 2023; Ahn et al., 2022, 2023; Zhu et al., 2023; Wang et al., 2023) and the catapult phenomenon (Lewkowycz et al., 2020; Kalra and Barkeshli, 2023; Meltzer and Liu, 2023; Zhu et al., 2024a, b). Compared to these works, our analysis extends to even larger (near-critical) step sizes. Another line of work shows how a large step size can enhance feature learning in one step of gradient descent (Ba et al., 2022; Dandi et al., 2024; Moniri et al., 2025), also comparing different parametrizations (Sonthalia et al., 2025). Similar observations about the role of the step size in feature learning have also been made for SGD (Andriushchenko et al., 2023; Lu et al., 2024) and pre-training (Sadrtdinov et al., 2024). Ziyin et al. (2022) observed that for a certain range of step sizes, SGD can have undesirable behavior, such as convergence to local maxima. For linear networks, Kreisler et al. (2023) identified a monotonically decreasing quantity (sharpness) along the gradient descent trajectories. Wang et al. (2022) showed large step size induces an implicit bias towards balanced minimizers in matrix factorization. Crăciun and Ghoshdastidar (2024) proved the existence of a step size threshold above which the algorithm diverges. A similar problem is studied by Marion and Chizat (2024). Large-step-size gradient descent has also been investigated in logistic regression (Wu and Su, 2023; Wu et al., 2024; Meng et al., 2024), and some of the analysis has been further extended to shallow networks (Cai et al., 2024).
@@
Van Den Doel and Ascher (2012) empirically observed chaos, specifically positive finite-time Lyapunov exponents, for several variants of steepest descent methods. The phenomenon named period-doubling bifurcation route to chaos has been widely observed in recent literature (Kong and Tao, 2020; Chen and Bruna, 2023; Chen et al., 2024b; Meng et al., 2024; Danovski et al., 2024; Ghosh et al., 2025). Among them, only Kong and Tao (2020); Chen et al. (2024b) provided rigorous analyses for the chaotic dynamics. They showed the emergence of Li-Yorke chaos, i.e., the existence of periodic orbits of arbitrary periods, for univariate training losses. In comparison, our setting is high-dimensional. Additionally, we establish not only the existence of all periodic orbits, but also the sensitivity of the limiting point to initialization, which is more relevant to practical optimization, particularly the implicit bias of the optimization algorithm.
@@ 2 Preliminaries
Gradient optimization in problem (1) has been extensively studied, especially in small-step-size regimes (see, e.g., Saxe et al., 2014; Arora et al., 2019; Yun et al., 2021; Du et al., 2018; Li et al., 2021; Min et al., 2023; Chen et al., 2024a). Its landscape enjoys a favorable structure: the global minimum is attained and every stationary point is either a global minimizer or a strict saddle (Li et al., 2019b; Valavi et al., 2020; Zhou et al., 2022). Nevertheless, this problem retains a key complexity typical of neural network optimization: global minimizers can differ substantially (although they yield the same end-to-end matrix). In particular, in the unregularized case, both the parameter norm $\|U\|_{F}^{2}+\|V\|_{F}^{2}$ and the imbalance $\|UU^{\top}-VV^{\top}\|_{F}^{2}$ can be arbitrarily large on the set of minimizers. This makes the problem a natural testbed for studying how hyperparameter choices affect the implicit biases of parameter optimization algorithms. We note that other forms of regularization have also been studied in the literature, such as $\|UU^{\top}-VV^{\top}\|_{F}^{2}$ (Tu et al., 2016; Ge et al., 2017).
@@
We define the gradient descent update map $\mathrm{GD}_{\eta}(\theta)=\theta-\eta\nabla L(\theta)$ so that $(U_{t+1},V_{t+1})=\mathrm{GD}_{\eta}(U_{t},V_{t})$. The basin of attraction of a stationary point $\theta^{*}$ of $L$ is the set of all initializations that converge to $\theta^{*}$, $\{\theta\colon\lim_{N\to\infty}\mathrm{GD}^{N}_{\eta}(\theta)=\theta^{*}\}$. [^1] In dynamical systems the basin of attraction is often defined for attractors. Here we extend the terminology to include all stationary points, such as saddles, for simplicity of presentation. The convergence region for step size $\eta$, $\mathcal{D}_{\eta}$, is the union of the basins of attraction of all global minimizers. The critical step size $\eta^{*}(\bar{U},\bar{V})$ for an initialization $(\bar{U},\bar{V})$ is defined as $\eta^{*}(\bar{U},\bar{V})=\sup\left\{\eta\colon\lim_{N\to\infty}\mathrm{GD}^{N}_{\eta}(\theta)\in\mathcal{M}\right\}$, i.e., the supremum of the step sizes that allow convergence to a global minimizer, where $\mathcal{M}=\left\{\theta\colon L(\theta)=\min_{\theta^{\prime}}L(\theta^{\prime})\right\}$ denotes the set of all global minimizers. A set $S$ is said to be invariant under $\mathrm{GD}_{\eta}$ if $\mathrm{GD}_{\eta}(S)\subset S$.
@@
We introduce notions for describing fractal geometry. A fractal is typically defined as a shape that exhibits self-similarity and fine structure at arbitrarily small scales. Formally, we say a set $S\subset\mathbb{R}^{n}$ is self-similar with degree k if there exist $k$ homeomorphisms, $\phi_{i}\colon S\to S,i=1,\cdots,k$, that satisfy (i) $S=\cup_{i=1}^{k}\phi_{i}(S)$ and (ii) there exists an open set $O\subset S$ such that $\cup_{i=1}^{k}\phi_{i}(O)\subset O$ and $(\phi_{i}(O))_{i=1}^{k}$ are pairwise disjoint. Condition (i) states that $S$ can be covered by $k$ smaller copies of itself. Condition (ii), which is known as the open set condition, ensures that those copies do not overlap much. This definition is closely related to an Iterated Function System (IFS), a standard tool for analyzing fractals (see, e.g., Hutchinson, 1981; Falconer, 2013). However, unlike IFS where the maps are required to be contractive and the set $S$ to be compact, the shapes considered in our study may be unbounded.
@@
Finally, we introduce notions related to chaos. Although there is no universal definition of chaos, one common characterization of chaos is the sensitivity to initialization, which is often known as the butterfly effect. In the context of optimization, this means that infinitesimal perturbations of the initial condition can lead to substantially different training outcomes (e.g., turning convergence into divergence, or shifting convergence from one minimizer to another qualitatively different minimizer). In dynamical systems, a key measure of chaos is the topological entropy. Informally, the topological entropy $h(F)$ of a dynamical system $F$ measures the exponential growth rate of the number of distinct trajectories of $F$ as a function of the trajectory length. We defer the formal definition to Appendix B.1. A positive topological entropy is widely regarded as a hallmark of chaos (see, e.g., Katok et al., 1995; Robinson, 1998; Vries, 2014). In this paper, we adopt the above notions for fractals and chaos. We note that different definitions and settings exist, and discuss their relation to our study in Appendix A.